A7 year old child has been prescribed hydroxyzine 10 mg/5 ml syrup at the following dose: 10 mg at night for the first week, then 20 mg at night for the second week, then 30mg at night for the third week, then 30 mg bd. What volume of syrup should be dispensed to provide 28 days supply for this patient?
Answer #1

ANSWER

* Volume of syrup dispensed to provide 28 days.

              Prescribed dose for first week = 10 mg * 7 days.

                                                          = 70 mg / week.

          Prescribed dose for second week = 20 mg * 7 days.

                                                                = 140 mg / week.

          Prescribed dose for third week = 30 mg * 7 days.

                                                            = 210 mg / week.

      Prescribed dose for last week   = 30 mg( bd) * 7 days.

                                                                 = (30 * 7 ) 2

                                                                = 420 mg / week.

       Total amount of syrup needed (mg) = 70 + 140 + 210 + 420

                                                                       = 840 mg.

         Volume of syrup needed for 28 days (ml)

                         Prescribed hydroxyzine dose = 10 mg /5 ml.

                                                                   1 ml = 2 mg.

                            So

                                Total volume of syrup needed ( ml)

                                                                        1 ml = 2 mg

                    ( Assume x as required ml)        x ml = 840 mg

                                                                      2x = 840 * 1

                                                                            x   = 840 / 2

                                                                                 = 420 ml.

                        * Volume of syrup should be dispensed is 420 ml.